OpenStudy (anonymous):

f(x)=1/(x^2+1) g(x)=2x^2-x+4 Evaluate: f(g(x) )=(f∘g)(x)

OpenStudy (anonymous):

When you see the expression f(g(x)), the g(x) is inside f's ( ), right? So...think of it as putting g(x) inside the f(x) function.

OpenStudy (anonymous):

How do you put g(x) inside f(x) = 1/(x^2 + 2) ? You simply replace that x that is in the denominator of the function with g(x)...that is, you replace the x with 2x^2-x+4

OpenStudy (anonymous):

f(x) = 1/(x^2+1)...................that's given. Now write f(x) replacing the x with 2x^2-x+4 f(g(x)) = 1/((2x^2-x-4)^2+1)

OpenStudy (anonymous):

Hello..............are you there?

OpenStudy (anonymous):

Sorry about that had to return to work. Thanks for all the help. When i working on this problem i was having trouble with how it should look being that f(x) is a rational.